Reasoning missing number question....
Q- 4,3 and 7 have a triangular relation which get 42 as outcome and similarly 6,3 and 8 have same format relation with each other which get 91 as outcome so find the outcome if 5,3 and 9 have the same order relation with each other? find missing number.
1
Expert's answer
2020-04-09T14:19:42-0400
"7^2-(4^2+3^2)=49-(16+9)=49-25=24"

If we rearrange the digits we obtain "42."



"8^2-(6^2+3^2)=64-(36+9)=64-45=19"

If we rearrange the digits we obtain "91."



"9^2-(5^2+3^2)=81-(25+9)=81-34=47"

If we rearrange the digits we obtain "74."


The missing number is "74."